# Token-Budget-Aware LLM Reasoning

### v4

@@ Abstract
Reasoning is critical for large language models (LLMs) to excel in a wide range of tasks. While methods like Chain-of-Thought (CoT) reasoning and enhance LLM performance by decomposing problems into intermediate steps, they also incur significant overhead in token usage, leading to increased costs. We find that the reasoning process of current LLMs is unnecessarily lengthy and it can be compressed by including a reasonable token budget in the prompt, but the choice of token budget plays a crucial role in the actual compression effectiveness. We then propose a token-budget-aware LLM reasoning framework that dynamically adjusts the number of reasoning tokens based on the reasoning complexity of each problem. Experiments show that our method effectively reduces token costs in CoT reasoning with only a slight performance reduction, offering a practical solution to balance efficiency and accuracy in LLM reasoning. Code: https://github.com/GeniusHTX/TALE.
@@
Tingxu Han [^1] Equal Contribution, Project Lead.1, Zhenting Wang [^1] Equal Contribution, Project Lead.2, Chunrong Fang [^2] Corresponding Author.1, Shiyu Zhao2, Shiqing Ma3, Zhenyu Chen1 1Nanjing University 2Rutgers University 3UMass Amherst
@@ 1 Introduction
Reasoning plays a crucial role in enabling large language models (LLM) to perform effectively across a wide range of tasks Zhou et al. (2022); Hao et al. (2023, 2024a). A variety of methods have been proposed to enhance the reasoning capabilities of large language models Suzgun et al. (2022); Wang et al. (2023); Feng et al. (2023); Xie et al. (2024). Among these, Chain-of-Thought (CoT) Wei et al. (2022) is the most representative and widely adopted approach. It enhances the reliability of the model’s answers by guiding large language models with the prompt “Let’s think step by step”, encouraging them to decompose the problem into intermediate steps and solve each before arriving at the final answer. Figure 1a and Figure 1b illustrate an intuitive example. Observe that without CoT, the LLM produces incorrect answers to the question. With a CoT-enhanced prompt, the LLM systematically breaks the question into multiple steps and reasons through each step sequentially. By addressing each step incrementally, the LLM eventually arrives at the correct answer. Recent reasoning models, such as OpenAI O1 OpenAI (2024c) and DeepSeek R1 Guo et al. (2025), integrate CoT into their design. Notably, these models can perform CoT reasoning even without explicit prompting.
@@
Although reasoning enhancement approaches such as CoT impressively improve LLM performance, they produce substantial additional overhead, specifically in the form of the increased number of tokens produced Wei et al. (2022); Feng et al. (2023); Yao et al. (2024a). As shown in Figure 1b, the answer to prompt with CoT has notably higher token costs due to the detailed intermediate reasoning steps included in the output. Such high token costs can lead to significant expenses, including increased computational resource usage and longer running times during the LLM inference, ultimately resulting in significant additional monetary and energy costs.
@@ 2 Related Work
Token Cost of LLM. Although the above methods enhance reasoning accuracy, they often increase token usages, posing challenges to efficiency Wang et al. (2024); Chiang and Lee (2024); Bhargava et al. (2023). Consequently, it is important to mitigate token consumption while maintaining the model performance. To address this issue, Li et al. (2021) introduces a multi-hop processing technique designed to filter out irrelevant reasoning. While effective, this approach is limited to traditional neural networks, such as PALM Bi et al. (2020), and lacks adaptability to large language models (LLMs). Zheng et al. (2024) aims to improve LLM inference speed by predicting response lengths and applying a scheduling algorithm to enhance efficiency. However, it is constrained to scheduling level, and it does not reduce the actual token costs. Hao et al. (2024b) reduces token usage by substituting decoded text tokens with continuous latent tokens. However, its application is currently restricted to small-scale, early language models like GPT-2 Radford et al. (2019). Additionally, it significantly impacts reasoning accuracy, resulting in over a 20% relative accuracy reduction on benchmarks such as GSM8K Cobbe et al. (2021).
@@ 4 Searching Optimal Token Budget
(a) GPT-4o-mini budget search.
(b) GPT-4o-mini token cost.
(c) Yi-lightning budget search.
(d) Yi-lightning token cost.
@@
Vanilla Method for Optimal Budget Search. An intuitive method is finding the minimal needed tokens as the budget, ensuring that the LLM can still produce correct and accurate responses within this constraint. To find the minimal token budget required for each question, we utilize a binary search-based minimal budget search algorithm. Before initiating the search process, we first apply the vanilla CoT to generate an answer for each question, as illustrated in Figure 1b. The number of tokens in the resulting answer is then calculated and designated as the right boundary for search, denoted by $right$. The function isFeasible is used to determine the feasibility of a budget. A budget is considered feasible here if the CoT prompt with that budget preserves the correctness of the answer. Algorithm 1 showcases the details. Given the feasibility function, large language model $\mathcal{M}$, question ${\bm{x}}$ and label $y$ as the input, Algorithm 1 first calculates the right boundary of search (line 2). With 00 as the left boundary, the current possible budget $\beta$ is computed as the midpoint of 00 and $right$ (line 3). We use $\beta_{0}$ to record the previously searched budget (line 4). While the current $\beta$ is feasible, the algorithm updates $\beta$ by recalculating the midpoint (line 7) and adjusts the search bounds accordingly to narrow the range (line 9). Once the loop ends, the final budget $\beta$ is returned as the searched result (line 12). Algorithm 1 is designed to find the minimal budget efficiently. However, we observe that the minimal budget required to produce a correct answer is not necessarily the optimal budget. When the budget is unreasonably small, the actual token cost often exceeds that of cases where a larger budget is used.
@@
Observation of Token Elasticity. During our minimal budget search process, we observe a “token elasticity” phenomenon as we approach the minimal budget. Specifically, as Algorithm 1 progresses, we aim to identify the minimal budget that still ensures the answer’s correctness. However, we find that if the budget is reduced beyond a certain range, the token cost increases, indicating that further reductions in the budget lead to increasing token consumption. Figure 2 showcases the evidence. The x-axis represents the iterations of the budget binary search, with the budget values decreasing progressively. The y-axis in Figure 2b and Figure 2d show the corresponding token costs at each budget search iteration. We can observe that when the searched budget crosses a range for reasonable budgets, the token costs begin to increase gradually. Figure 1c also shows an example. As observed, when a small token budget (e.g., 10 tokens) is used, the real token cost is significantly higher compared to scenarios where a reasonable token budget is allocated (i.e., Figure 1d).
@@ 5.2 Estimation and Prompting (TALE-EP)
Our observations on token elasticity (Section 4) indicate that only a well-chosen budget within a reasonable range can effectively minimize token costs while preserving LLM performance. The optimal budget, found using Algorithm 1 and Algorithm 2, lies within this range and achieves a satisfying trade-off between efficiency and performance. Building on this insight, we introduce a token budget aware reasoning method by zero-shot-based token budget estimation and prompting the reasoning LLM. TALE-EP leverages the reasoning capabilities of the LLM as an estimator. Figure 4 provides an overview of TALE-EP’s workflow. The goal of TALE-EP is to construct a token-budget-aware prompt that maintains performance comparable to vanilla CoT while reducing token costs. To achieve this balance, TALE-EP follows a two-phase approach: budget estimation and prompt construction. Given a question, TALE-EP first estimates a reasonable token budget that closely aligns with the optimal searched budget. By default, we use the reasoning LLM itself with a zero-shot estimation prompt as the budget estimator. Figure 5 demonstrates the budget estimation prompt. Using this estimate, it then crafts a token-budget-aware prompt and feeds it into the LLM to generate the final answer. Figure 7c illustrates this process with a concrete example. The key intuition behind TALE-EP is inspired by human-like thinking. When solving a mathematical problem, a person may take time to compute the exact answer but can quickly estimate the effort required to solve it. For instance, when comparing a primary school arithmetic question to a college-level calculus problem, one may not immediately provide the solutions but can easily infer that the former takes only seconds while the latter requires significantly more time. Section A.2 evaluates the effectiveness of our budget estimation approach, demonstrating that the budgets estimated by advanced LLMs (e.g., GPT-4o-mini) are generally close to the optimal searched budget and deliver competitive performance.
@@ 6 Evaluation
In this section, we provide the experiment results to evaluate the effectiveness of two versions of TALE, TALE-EP and TALE-PT.
@@ 6.2 Effectiveness of TALE-EP
Table 2 compares TALE-EP with other prompt engineering methods across seven datasets, evaluating accuracy, output tokens, and expenses. Effective prompts should maximize accuracy while minimizing token usage and cost. Direct Answering is the most cost-efficient (14.57 tokens, 25.37 expense) but with low accuracy (52.31%). Vanilla CoT achieves the highest accuracy (83.75%) but at a high cost (461.25 tokens, 289.78 expense). TALE-EP balances performance and efficiency, achieving 81.03% accuracy while reducing token usage to 32% and expenses to 41% of Vanilla CoT. On GSM8K, it even surpasses Vanilla CoT with 84.46% accuracy. Note that expense is not directly proportional to output tokens because it also accounts for input and cached tokens. TALE-EP reduces token costs by 68.64% on average, offering a scalable, cost-effective solution for budget-constrained reasoning tasks.
To further evaluate the generalization of TALE-EP across different LLMs. We conduct experiments across Yi-lightning, GPT-4o-mini, GPT-4o and o3-mini on MathBench-College. Table 3 illustrates the results, showing TALE-EP’s ability to reduce output tokens and expenses while maintaining competitive accuracy significantly. TALE-EP achieves substantial token savings, reducing output tokens by 64.63% on average, compared to Vanilla CoT. Expense reductions are equally notable, with costs decreasing by 45.30% on average. Despite these cost savings, TALE-EP maintains strong accuracy, achieving 76.67% on Yi-lightning, 70.00% on GPT-4o-mini, and 80.00% on GPT-4o, comparable to Vanilla CoT. These results highlight TALE-EP’s effectiveness in balancing cost efficiency and reasoning performance across diverse LLM architectures. The observed accuracy drop is most significant for GPT-4o-mini. This could be attributed to its smaller number of parameters, which makes it more challenging to answer correctly within a limited response reasoning length.
@@ A.1 Definition of Ideal Budget Range
Intuitively, a higher in-range accuracy and a lower out-range distance indicate a better estimated budget. During our evaluation, the in-range accuracy is 60.61%, and the out-of-range distance is 109.64. It indicates that more than two-thirds of estimated budgets are located in the ideal range. For those out-of-range samples, they have an offset of 109.64 tokens on average.

### v5

@@ Abstract
Reasoning is critical for large language models (LLMs) to excel in a wide range of tasks. While methods like Chain-of-Thought (CoT) reasoning and enhance LLM performance by decomposing problems into intermediate steps, they also incur significant overhead in token usage, leading to increased costs. We find that the reasoning process of current LLMs is unnecessarily lengthy and it can be compressed by including a reasonable token budget in the prompt, but the choice of token budget plays a crucial role in the actual compression effectiveness. We then propose a token-budget-aware LLM reasoning framework that dynamically adjusts the number of reasoning tokens based on the reasoning complexity of each problem. Experiments show that our method effectively reduces token costs in CoT reasoning with only a slight performance reduction, offering a practical solution to balance efficiency and accuracy in LLM reasoning. Code: https://github.com/GeniusHTX/TALE [^1] Also available at https://www.gitlink.org.cn/txhan/TALE.
@@
Tingxu Han [^1] Equal Contribution.1, Zhenting Wang [^1] Equal Contribution. [^2] Start the project and propose the idea.2, Chunrong Fang [^3] Corresponding Author.1, Shiyu Zhao2, Shiqing Ma3, Zhenyu Chen1 1Nanjing University 2Rutgers University 3UMass Amherst
@@ 1 Introduction
Reasoning plays a crucial role in enabling large language models (LLM) to perform effectively across a wide range of tasks Zhou et al. (2022); Hao et al. (2023, 2024a); Jin et al. (2024a); Wang et al. (2024b, 2025). A variety of methods have been proposed to enhance the reasoning capabilities of large language models Suzgun et al. (2022); Wang et al. (2023); Feng et al. (2023); Xie et al. (2024). Among these, Chain-of-Thought (CoT) Wei et al. (2022) is the most representative and widely adopted approach. It enhances the reliability of the model’s answers by guiding large language models with the prompt “Let’s think step by step”, encouraging them to decompose the problem into intermediate steps and solve each before arriving at the final answer. Figure 1a and Figure 1b illustrate an intuitive example. Observe that without CoT, the LLM produces incorrect answers to the question. With a CoT-enhanced prompt, the LLM systematically breaks the question into multiple steps and reasons through each step sequentially. By addressing each step incrementally, the LLM eventually arrives at the correct answer. Recent reasoning models, such as OpenAI O1 OpenAI (2024c) and DeepSeek R1 Guo et al. (2025), integrate CoT into their design. Notably, these models can perform CoT reasoning even without explicit prompting.
@@
Although reasoning enhancement approaches such as CoT impressively improve LLM performance, they produce substantial additional overhead, specifically in the form of the increased number of tokens produced Wei et al. (2022); Feng et al. (2023); Yao et al. (2024a); Jin et al. (2024b). As shown in Figure 1b, the answer to prompt with CoT has notably higher token costs due to the detailed intermediate reasoning steps included in the output. Such high token costs can lead to significant expenses, including increased computational resource usage and longer running times during the LLM inference, ultimately resulting in significant additional monetary and energy costs.
@@ 2 Related Work
Token Cost of LLM. Although the above methods enhance reasoning accuracy, they often increase token usages, posing challenges to efficiency Wang et al. (2024a); Chiang and Lee (2024); Bhargava et al. (2023). Consequently, it is important to mitigate token consumption while maintaining the model performance. To address this issue, Li et al. (2021) introduces a multi-hop processing technique designed to filter out irrelevant reasoning. While effective, this approach is limited to traditional neural networks, such as PALM Bi et al. (2020), and lacks adaptability to large language models (LLMs). Speculative decoding Leviathan et al. (2023) aims to accelerate decoding by generating drafts using smaller models and verifying them with larger models, which is over-dependent on the alternative small approximation model. LLM routing Ding et al. (2024) queries to different LLMs based on quality-cost trade-offs, but it cannot reduce the token usage on the specific LLM for a given query. Zheng et al. (2024) aims to improve LLM inference speed by predicting response lengths and applying a scheduling algorithm to enhance efficiency. However, it is constrained to scheduling level, and it does not reduce the actual token costs. Hao et al. (2024b) reduces token usage by substituting decoded text tokens with continuous latent tokens. However, its application is currently restricted to small-scale, early language models like GPT-2 Radford et al. (2019). Additionally, it significantly impacts reasoning accuracy, resulting in over a 20% relative accuracy reduction on benchmarks such as GSM8K Cobbe et al. (2021).
@@ 4 Searching Optimal Token Budget
(a) GPT-4o-mini budget search.
(b) GPT-4o-mini token cost.
(c) Yi-lightning budget search.
(d) Yi-lightning token cost.
Vanilla Method for Optimal Budget Search. An intuitive method is finding the minimal needed tokens as the budget, ensuring that the LLM can still produce correct and accurate responses within this constraint. The goal of the search algorithm is not to directly determine task difficulty, but to identify the minimum token budget under which the model can still produce a correct answer. Specifically, the algorithm conducts a binary search over different token budgets and selects the shortest budget that maintains correctness. This approximates the minimum reasoning length required for the model to solve the given problem. To find the minimal token budget, we first propose an “implicit monotonicity assumption” that when the model outputs a wrong prediction at a budget value, it always predicts incorrectly when under this budget value, and when the model outputs a correct prediction at a budget value, it always predicts correctly when above this budget value. To empirically assess the validity of this assumption, we conduct an additional analysis. Specifically, we define a sample as monotonic if all predictions above the optimal budget are correct, and all predictions below it are incorrect. We randomly sample from the GSM8K dataset as test data and find that 90.91% of the samples satisfy this monotonicity condition. This suggests that while the assumption may not strictly hold in every instance, it is a reasonable and effective approximation in practice for guiding budget search. An intuitive monotonic sample is illustrated as Table 2. Based on the “implicit monotonicity assumption”, we further design a binary search-based minimal budget search algorithm detailed in Algorithm 1.
Table 2: An intuitive monotonic example. $\beta^{*}$ is the searched optimal budget. The budget row displays scaled budgets ranging from $2^{-2}$ to $2^{2}\cdot\beta^{*}$.
Budget($*\beta^{*}$) $2^{-2}$ $2^{-1}$ $1$ $2^{1}$ $2^{2}$
Prediction False False True True True
Before initiating the search process, we first apply the vanilla CoT to generate an answer for each question, as illustrated in Figure 1b. The number of tokens in the resulting answer is then calculated and designated as the right boundary for search, denoted by $right$. The function isFeasible is used to determine the feasibility of a budget. A budget is considered feasible here if the CoT prompt with that budget preserves the correctness of the answer. Algorithm 1 showcases the details. Given the feasibility function, large language model $\mathcal{M}$, question ${\bm{x}}$ and label $y$ as the input, Algorithm 1 first calculates the right boundary of search (line 2). With 00 as the left boundary, the current possible budget $\beta$ is computed as the midpoint of 00 and $right$ (line 3). We use $\beta_{0}$ to record the previously searched budget (line 4). While the current $\beta$ is feasible, the algorithm updates $\beta$ by recalculating the midpoint (line 7) and adjusts the search bounds accordingly to narrow the range (line 9). Once the loop ends, the final budget $\beta$ is returned as the searched result (line 12). Algorithm 1 is designed to find the minimal budget efficiently. However, we observe that the minimal budget required to produce a correct answer is not necessarily the optimal budget. When the budget is unreasonably small, the actual token cost often exceeds that of cases where a larger budget is used. We also further formalize the optimal budget search process detailed in Section A.6.
@@
Observation of Token Elasticity. During our minimal budget search process, we observe a “token elasticity” phenomenon as we approach the minimal budget. Specifically, as Algorithm 1 progresses, we aim to identify the minimal budget that still ensures the answer’s correctness. However, we find that if the budget is reduced beyond a certain range, the token cost increases, indicating that further reductions in the budget lead to increasing token consumption. Figure 2 showcases the evidence. The x-axis represents the iterations of the budget binary search, with the budget values decreasing progressively. The y-axis in Figure 2b and Figure 2d show the corresponding token costs at each budget search iteration. When the searched budget drops below a reasonable range, the token cost increases. This happens because the model, unable to meet the tight constraint, ignores it and reverts to longer reasoning. In other words, the model effectively “gives up” on complying with the instruction, resulting in longer outputs and redundant token costs. This explains the non-monotonicity observed in Figure 2: token usage initially decreases as the budget tightens but eventually rebounds when the budget becomes too small to be feasible. We will make this intuition clearer in the revised version. Figure 1c also shows an example. As observed, when a small token budget (e.g., 10 tokens) is used, the real token cost is significantly higher compared to scenarios where a reasonable token budget is allocated (i.e., Figure 1d).
@@
Overhead of Budget Search. For the overhead, note that the budget search mainly serves to reveal token compression potential, motivating TALE’s design. For TALE-PT, the budget search is used once offline to generate training targets for post-training. Since this is a one-time pre-processing step, it does not incur additional costs during actual model deployment. For TALE-EP, we clarify that the budget search is not needed at all. TALE-EP relies on a lightweight, zero-shot budget estimator that directly predicts a reasonable token budget without any iterative search, making it training-free and highly efficient at inference time. To quantify the budget search overhead for TALE-PT data generation, we measured the total time needed to run the search algorithm and produce the optimal budget dataset. On GSM8K (7473 samples), this process takes approximately 354 minutes on an A100 GPU, which we consider acceptable given that it is a one-time offline cost for training.
@@ 5.2 Estimation and Prompting (TALE-EP)
Our observations on token elasticity (Section 4) indicate that only a well-chosen budget within a reasonable range can effectively minimize token costs while preserving LLM performance. The optimal budget, found using Algorithm 1 and Algorithm 2, lies within this range and achieves a satisfying trade-off between efficiency and performance. Building on this insight, we introduce a token budget aware reasoning method by zero-shot-based token budget estimation and prompting the reasoning LLM. TALE-EP leverages the reasoning capabilities of the LLM as an estimator. Figure 4 provides an overview of TALE-EP’s workflow. The goal of TALE-EP is to construct a token-budget-aware prompt that maintains performance comparable to vanilla CoT while reducing token costs. To achieve this balance, TALE-EP follows a two-phase approach: budget estimation and prompt construction. Given a question, TALE-EP first estimates a reasonable token budget that closely aligns with the optimal searched budget. By default, we use the reasoning LLM itself with a zero-shot estimation prompt as the budget estimator. Figure 5 demonstrates the budget estimation prompt, which will guide the model to evaluate the question as a whole.. Using this estimate, it then crafts a token-budget-aware prompt and feeds it into the LLM to generate the final answer. Figure 7c illustrates this process with a concrete example. The key intuition behind TALE-EP is inspired by human-like thinking. When solving a mathematical problem, a person may take time to compute the exact answer but can quickly estimate the effort required to solve it. For instance, when comparing a primary school arithmetic question to a college-level calculus problem, one may not immediately provide the solutions but can easily infer that the former takes only seconds while the latter requires significantly more time. Section A.2 evaluates the effectiveness of our budget estimation approach, demonstrating that the budgets estimated by advanced LLMs (e.g., GPT-4o-mini) are generally close to the optimal searched budget and deliver competitive performance.
@@ 6 Evaluation
In this section, we provide the experiment results to evaluate the effectiveness of two versions of TALE, TALE-EP and TALE-PT. The comparisons of the two implementations are detailed in Section A.6.
@@ 6.2 Effectiveness of TALE-EP
Table 3 compares TALE-EP with other prompt engineering methods across seven datasets, evaluating accuracy, output tokens, and expenses. Effective prompts should maximize accuracy while minimizing token usage and cost. Direct Answering is the most cost-efficient (14.57 tokens, 25.37 expense) but with low accuracy (52.31%). Vanilla CoT achieves the highest accuracy (83.75%) but at a high cost (461.25 tokens, 289.78 expense). TALE-EP balances performance and efficiency, achieving 81.03% accuracy while reducing token usage to 32% and expenses to 41% of Vanilla CoT. On GSM8K, it even surpasses Vanilla CoT with 84.46% accuracy. Note that expense is not directly proportional to output tokens because it also accounts for input and cached tokens. TALE-EP reduces token costs by 68.64% on average, offering a scalable, cost-effective solution for budget-constrained reasoning tasks. For resource-rich scenarios, we evaluate TALE-EP under a larger token budget as detailed in Section A.8.
To further evaluate the generalization of TALE-EP across different LLMs. We conduct experiments across Yi-lightning, GPT-4o-mini, GPT-4o and o3-mini on MathBench-College. Table 4 illustrates the results, showing TALE-EP’s ability to reduce output tokens and expenses while maintaining competitive accuracy significantly. TALE-EP achieves substantial token savings, reducing output tokens by 64.63% on average, compared to Vanilla CoT. Expense reductions are equally notable, with costs decreasing by 45.30% on average. Despite these cost savings, TALE-EP maintains strong accuracy, achieving 76.67% on Yi-lightning, 70.00% on GPT-4o-mini, and 80.00% on GPT-4o, comparable to Vanilla CoT. These results highlight TALE-EP’s effectiveness in balancing cost efficiency and reasoning performance across diverse LLM architectures. The observed accuracy drop is most significant for GPT-4o-mini. This could be attributed to its smaller number of parameters, which makes it more challenging to answer correctly within a limited response reasoning length. We also evaluate the applicability of TALE on more tasks, the results are illustrated in Section A.5. The efficiency analysis of TALE-EP is in Section A.7
@@ A.1 Definition of Ideal Budget Range
Intuitively, a higher in-range accuracy and a lower out-range distance indicate a better estimated budget. During our evaluation, the in-range accuracy is 60.61%, and the out-of-range distance is 109.64. It indicates that more than two-thirds of estimated budgets are located in the ideal range. For those out-of-range samples, they have an offset of 109.64 tokens on average. Figure 9 illustrates the successful and failed estimated cases intuitively. The prompt we use for budget estimation is as follows:
‘‘Task: Analyze the given question and estimate the minimum number of tokens required for reasoning.’’
This prompt encourages the model to evaluate the question as a whole, including reasoning depth, structure, completeness, and surface-level difficulty.
@@ A.3 Details of TALE’s Implementation
(a) Successful case.
(b) Failed case.
Figure 9: An intuitive example for successful and failed cases of prompt budget estimation in TALE-EP.
A.4 Comparison of TALE-EP and TALE-PT.
In this section, we compare the performance of TALE-EP and TALE-PT (including SFT and DPO). Specifically, we utilize Llama-3.1-8B-Instruct as both the budget estimator for TALE-EP and the base model for TALE-PT, evaluated on GSM8K. Table 6 illustrates the evidence. As the zero-shot-based estimator tends to predict relatively low budgets, TALE-EP achieves lower token usage but at the cost of slightly reduced accuracy. In contrast, both variants of TALE-PT achieve higher accuracy with more tokens, as their training data is constructed using optimal budget search, which enforces answer correctness as a strict constraint even with a higher token costs. This highlights a trade-off between strict correctness preservation (in TALE-PT) and token efficiency (in TALE-EP).
Table 6: Comparison of TALE-EP and TALE-PT.
Metrics TALE-EP TALE-PT
SFT DPO
ACC 71.82 78.57 74.11
Output Tokens 112.21 139.63 149.93
A.5 Applicability of TALE on More Tasks.
To further evaluate the applicability, we deploy TALE-EP on three additional open-ended generative tasks. As these tasks involve open-ended text generation, we adopt the BLEU metric Papineni et al. (2002) to quantify the similarity between generated outputs and reference texts. The results in Table 4 demonstrate that TALE-EP achieves comparable or even better BLEU scores than Vanilla CoT while using only around 40% of the output tokens, validating its effectiveness and applicability in broader generative tasks.
[…]
Tasks TALE-EP Vanilla CoT
BLEU Output Tokens BLEU Output Tokens
CS 0.07 44.39 0.2 134.05
ERG 0.005 60.34 0.006 175.37
CG 0.24 171.08 0.267 461.77
A.6 Formalizing the Budget Search.
For a given input ${\bm{x}}$, we define the search space as:
=$\mathcal{B}=\{\beta\in\mathbb{Z}^{+}\mid 0<\beta\leq T_{\text{vanilla}}(x)\}$
where $T_{\text{vanilla}}(x)$ is the number of tokens generated by Vanilla CoT. A candidate budget $\beta\in\mathcal{B}$ is considered feasible if:
$LLM(x,\beta)=y\text{ and }T(x,\beta)<T(x,\beta_{0})$
where $y$ is the ground-truth answer, $T(x,\beta)$ is the actual number of output tokens when answering $x$ under budget $\beta$, $\beta_{0}$ is the previously searched (larger) feasible budget. Our goal is to find:
$\beta^{*}=\arg\min_{\beta\in\mathcal{B}}T(x,\beta)\text{, subject to }LLM(x,\beta)=y$
To efficiently find $\beta^{*}$, we employ a binary search procedure guided by the feasibility function above, as detailed in Algorithm 1 and Algorithm 2.
A.7 Efficiency of TALE-EP.
Since TALE-EP requires one additional query, we further evaluate its end-to-end latency in this section. Specifically, we query the Llama-3.1-8B-Instruct model over the GSM8K-Zero dataset and measure both accuracy and average time cost. T he budget estimation query of TALE-EP is also considered. Although TALE-EP requires one additional query compared to Vanilla CoT, it is significantly more efficient, taking only 2.3 seconds per sample, while Vanilla CoT takes 10.2 seconds. This is because the primary factor influencing inference time is the number of output tokens, which TALE-EP effectively reduces.
A.8 Effectiveness of Larger Token Budget.
In scenarios with ample computational resources, the token budget could be larger for better performance. We simulate such a scenario by scaling the estimated budget by a factor $\alpha$ ($\alpha*\text{budget}$). As shown in the table below, increasing $\alpha$ from 1 to 2 leads to higher accuracy (from 67.33% to 72.66%) at the cost of more tokens (from 210.97 to 279.78), demonstrating that TALE-EP can flexibly adapt to different resource scenarios.
Table 8: The empirical evidence for “implicit monotonicity assumption”. $\bar{\beta}$ is the budget upper bound, which is the token cost of vanilla CoT. The budget row displays scaled budgets ranging from $2^{-2}$ to $2^{2}\cdot\beta^{*}$.
Budget($*\beta^{*}$) $2^{-5}$ $2^{-4}$ $2^{-2}$ $2^{0}$
ACC 69.23 75.82 75.82 76.92
Output Tokens 222.69 222.42 244.61 653.53
A.9 Empirical Evidence for the “Implicit Monotonicity Assumption”.
In this section, we give empirical evidence to support the “implicit monotonicity assumption” for our search algorithm. Specifically, for a set of questions, we vary the budget across a range of values (e.g., $2^{-5},2^{-4},2^{-2},2^{0}$ times budget upper bound, which is the token cost of vanilla CoT), and record the corresponding accuracy and average output tokens at each point. Table 8 illustrates the empirical evidence. Observe that it roughly follows from the monotonicity property. The results demonstrate a consistent trend: accuracy generally increases or plateaus with increasing budgets, confirming a soft monotonicity in most cases.